A 4 kg block moving right at 6 m/s collides elastically with a 2 kg block moving right at 3 m/s. Find the first
(V1) velocity.
1
Expert's answer
2020-11-23T10:28:46-0500
"m_1v_1+m_2v_2=m_1v_1'+m_2v_2'\\0.5m_1v_1^2+0.5m_2v_2^2=0.5m_1v_1'^2+0.5m_2v_2'^2\\v_1'=\frac{m_1-m_2}{m_1+m_2}v_1+\frac{2m_2}{m_1+m_2}v_2\\v_1'=\frac{4-2}{4+2}6+\frac{2(2)}{4+2}3=4\frac{m}{s}"